OpenStudy (anonymous):

I am trying to integrate 2(tanx)(sec^2x-1) from 0 to pi/4

OpenStudy (anonymous):

So far I have distributed and broken up the integral into 2

terenzreignz (terenzreignz):

Interesting... let's see if we can play with trigonometry to simplify things a bit... \[\Large \int\limits_{0}^{\frac \pi 4}2\tan(x)[\sec^2(x)-1]dx\]

OpenStudy (anonymous):

I began by taking out the constant, then distributing the tanx in and splitting up the problem

terenzreignz (terenzreignz):

That's good. Do you know the integral of tan(x) ?

OpenStudy (anonymous):

\[-\ln \left| cosx \right|\]

terenzreignz (terenzreignz):

or \[\Large \ln|\sec(x)|\] Okay, so we have no problem :)

terenzreignz (terenzreignz):

\[\large \int\limits_{0}^{\frac \pi 4}2\tan(x)[\sec^2(x)-1]dx=\int\limits_{0}^{\frac \pi 4}2\tan(x)\sec^2(x)dx -\int\limits_{0}^{\frac \pi 4}2\tan(x)dx\]

terenzreignz (terenzreignz):

So, can you do it from here?

OpenStudy (anonymous):

I got \[\tan ^{2}x \] for the first integral

terenzreignz (terenzreignz):

What? NO! It should be \(\large \sec^2(x)\)

terenzreignz (terenzreignz):

LOL Just messing with you :) It could actually be either :D

terenzreignz (terenzreignz):

\[\Large \sec^2(x) = \tan^2(x) + 1\]

OpenStudy (anonymous):

lol, cool

terenzreignz (terenzreignz):

Anyway, so... moving on?

OpenStudy (anonymous):

I think I'm just having issues evaluating then :/

terenzreignz (terenzreignz):

Okay, in the end we have \[\LARGE \left.\tan^2(x)- 2\ln|\sec(x)|\right]_0^{\frac\pi4}\]

terenzreignz (terenzreignz):

Or if it makes things simpler, one tiny tweak with the properties of logs... \[\LARGE \left.\tan^2(x)- \ln|\sec^{\color{red}2}(x)|\right]_0^{\frac\pi4}\]

OpenStudy (anonymous):

nice, good tip

terenzreignz (terenzreignz):

Okay, so evaluating this... \[\Large \tan^2\left(\frac\pi4\right)-\ln\left|\sec^2\left(\frac\pi4\right)\right|=1-\ln2\]

terenzreignz (terenzreignz):

\[\Large \tan(0) - \ln |\sec^2(0)|=0-\ln(1)=0\]

terenzreignz (terenzreignz):

Et voila ^_^

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

I understand now

OpenStudy (anonymous):

I need to review those special angles

OpenStudy (anonymous):

thank you :)

terenzreignz (terenzreignz):

^_^ That's good